[Brahms-l] Shift report 20060611 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 11 2006 - 08:07:07 EDT
Shifter leader was rk
Shifters were te


RHIC had problems with timing in Blue, which weren't solved for 5 hours. We had beam from 2am to 2.40am, another injection at 6.30am, still taking data.
At the beginning of the shift problems with BBR1 mainframe, after reset works fine.
As we turned on the voltages to take data in the second injection today, BADTRIP showed voltage deviation in T3. On opening we discovered T3A1XU was 5250V!!! instead of 1200V. Turning it off and on don't work, we left it like this, the current looks reasonable (1.3muA) and the planes don't have more hits than the neighbouring.
Will try to contact Pawel Staszel.
Due to the above problems only 3 runs (16048-16052) were taken, all with FS@3@1723A, MRS@45@700B.
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 11 08:07:49 2006

This archive was generated by hypermail 2.1.8 : Sun Jun 11 2006 - 08:08:02 EDT